3. A rectangle has a length 6 more than its width. If the perimeter of the rectangle is 40 m, what are the length and width of the rectangle?


1
Expert's answer
2021-09-01T19:50:16-0400

Let the width of the rectangle be "(x) m"

Thus the length of the rectangle will be "(x+6)m"

Perimeter of a rectangle ="2(L+W)=2(x+x+6)cm"

"40=4x+12"

"4x=40-12"

"4x=28"

"x=7m"

Thus :Width =7m,Length =13m